# Aux-Think: Exploring Reasoning Strategies for Data-Efficient Vision-Language Navigation

### v3

@@ Abstract
Vision-Language Navigation (VLN) is a critical task for developing embodied agents that can follow natural language instructions to navigate in complex real-world environments. Recent advances in VLN by large pretrained models have significantly improved generalization and instruction grounding compared to traditional approaches. However, the role of reasoning strategies in navigation—an action-centric, long-horizon task—remains underexplored, despite Chain-of-Thought (CoT) reasoning’s demonstrated success in static tasks like visual question answering. To address this gap, we conduct the first systematic evaluation of reasoning strategies for VLN, including No-Think (direct action prediction), Pre-Think (reason before action), and Post-Think (reason after action). Surprisingly, our findings reveal the Inference-time Reasoning Collaps issue, where inference-time reasoning degrades navigation accuracy, highlighting the challenges of integrating reasoning into VLN. Based on this insight, we propose Aux-Think, a framework that trains models to internalize structured reasoning patterns through CoT supervision, while inferring action directly without reasoning in online prediction. To support this framework, we release R2R-CoT-320k, the first Chain-of-Thought annotated dataset for VLN. Extensive experiments show that Aux-Think reduces training effort greatly and achieves the best performance on the navigation success rate.
@@ 1 Introduction
Vision-and-Language Navigation (VLN) wu2024vision ; anderson2018vision ; gu2022vision represents a groundbreaking step toward enabling robots to understand natural language instructions and navigate complex, unfamiliar environments. As a foundational capability for embodied AI systems, VLN bridges the gap between perception and action, empowering robots to seamlessly interact with the real world. In particular, Vision-and-Language Navigation in continuous environments (VLN-CE) hong2022bridging ; wang2023gridmm ; an2022bevbert has emerged as a critical research focus, pushing the boundaries of autonomy and adaptability in dynamic, real-world scenarios.
@@
Traditional Vision-and-Language Navigation methods often rely on waypoint predictors krantz2021waypoint ; hong2022bridging ; krantz2022sim or topology maps zubair2021sasra ; chen2021topological ; chen2022weakly but struggle with generalization and the sim-to-real gap. With the advancements in Large Language Models (LLMs) yang2024qwen2 ; touvron2023llama and Vision-Language Models (VLMs) bai2025qwen2 ; liu2024nvila ; lin2023video , recent studies zhang2024navid ; zhang2024uni ; cheng2024navila shift toward action prediction via supervised fine-tuning on paired videos and instructions. Despite these advancements, most efforts emphasize training strategies zhu2020vision ; zeng2024poliformer , data organization han2024roomtour3d , or model architecture zheng2024towards for VLN. Chain-of-Thought (CoT), which explicitly generates intermediate reasoning before producing final answers wei2022chain , has shown success in enhancing reasoning capabilities across various LLM- and VLM-driven tasks, like video understanding feng2025video and tool usage lu2025ui . However, its application to VLN remains unexplored.
@@
Motivated by this gap, we present the first systematic study of reasoning strategies in VLN, comparing three paradigms: (1) No-Think, which predicts actions without explicit reasoning; (2) Pre-Think zhang2024chain , which performs CoT before action selection; and (3) Post-Think li2025think , which reasons after action prediction. Our key findings reveal a phenomenon we term “Inference-time Reasoning Collapse” (IRC): Introducing CoT via Pre-Think or Post-Think consistently harms navigation performance (Fig. 1(a)). CoT errors or hallucinations during inference result in incorrect actions, as shown in Fig. 2. We attribute this to a training–inference mismatch: CoT is only learned on optimal (oracle) trajectories, but for inference, agents often enter unfamiliar, off-distribution states where reasoning fails, leading to error accumulation along the trajectory and cascading navigation failures. This phenomenon highlights a fundamental limitation of multi-turn explicit reasoning in dynamic, partially observable environments, unlike the single-turn static tasks such as VQA or image captioning antol2015vqa ; wang2024videocot .
@@
To address the IRC issue, we propose Aux-Think, inspired by the dual-process theory of human learning evans2003two : during training, humans often rely on explicit reasoning to understand principles, but during execution, they focus on actions without consciously recalling those principles, much like a driver who no longer recites traffic rules while driving. Similarly, Aux-Think uses CoT as an auxiliary signal during training to guide the model in internalizing reasoning patterns. At inference time, the model no longer generates explicit reasoning, but instead directly predicts actions based on the internalized reasoning learned during training. This separation between learning and execution improves decision focus, reduces inference overhead and hallucinations, and leads to more accurate and stable navigation.
@@
Specially, in Aux-Think, the generation of CoT reasoning and navigation actions is decoupled into two distinct tasks during training: (1) generating navigation actions based on stepwise observations and language instructions as the primary task, and (2) generating the reasoning process for each step as an auxiliary task. By leveraging prompts to switch between these tasks, we independently supervise navigation actions and reasoning processes, effectively avoiding the negative interference caused by jointly training both tasks. During inference, Aux-Think directly predicts actions without intermediate reasoning, thereby eliminating the risk of errors introduced by CoT hallucinations.
@@
  • New Finding: We conduct a systematic comparison of reasoning strategies in VLN and reveal that inference-time reasoning, including Pre-Think and Post-Think, consistently underperforms direct action prediction (No-Think), termed the IRC issue. To our knowledge, this is the first exploration of CoT strategies on the VLN-CE task.
  • New Method: We propose Aux-Think, a novel training paradigm that uses CoT as auxiliary supervision while maintaining No-Think inference, achieving superior performance over other reasoning methods. Aux-Think pioneers a new perspective on CoT utilization and achieves the best performance on the navigation success rate.
@@
Figure 2: Illustration of CoT and Action Prediction Results Using Different Reasoning Strategies. Pre-Think generates incorrect actions (e.g., “move forward 75cm”) due to flawed CoT reasoning, such as “we have not crossed the room,” leading to significant trajectory deviation. Post-Think, which builds on Pre-Think’s output, inherits similar reasoning errors (e.g., “no obvious door or archway”) and makes the same wrong prediction. In contrast, Aux-Think correctly predicts “turn right 15 degrees” and follows a trajectory aligned with the ground truth. While Aux-Think does not rely on CoT during inference, it can optionally produce CoT via prompt switching—yet its action prediction remains accurate even when the generated CoT is of moderate quality. This highlights Aux-Think’s robustness to imperfect reasoning and its superior reliability in action prediction.
@@ 2.2 VLN with Large Pretrained Models
Recent advancements have seen the integration of large pre-trained models liu2024nvila ; bai2025qwen2 ; yang2024qwen2 , into VLN tasks. Early explorations of LLM in the VLN field usually use off-the-shelf large language models to select landmarks or waypoints in a zero-shot manner long2024discuss ; long2024instructnav ; shah2023lm ; zhou2024navgpt . Recent works have focused on fine-tuning the VLM to obtain the navigational Vision-Language-Action model. Notably, Poliformer zeng2024poliformer and NaVid zhang2024navid introduce a video-based monocular VLN, demonstrating navigation capabilities using monocular RGB videos without maps or depth input. Uni-NaVid zhang2024uni unifies various navigation tasks, including VLN, ObjectNav chaplot2020object , Embodied Question Answering das2018embodied , and Human-following islam2019person ; puig2023habitat , into a single model trained on a diverse dataset. NaVILA cheng2024navila further extends this approach by integrating VLN with legged robot locomotion skills in complex environments.
While these models have improved the alignment among visual understanding, language instructions, and navigation actions, they predominantly employ No-Think inference strategies, lacking reasoning mechanisms. Moreover, their performance gains often stem from leveraging extensive datasets, whereas our approach focuses on exploring reasoning strategies.
@@ 2.3 Reasoning Models
In our work, we conduct the first systematic comparison of Pre-Think, Post-Think, and No-Think reasoning strategies for VLN. Based on our findings, we propose Aux-Think, a novel framework that leverages CoT reasoning as auxiliary supervision during training while maintaining No-Think inference, thereby enhancing data efficiency and performance in VLN.
@@ 3.1 Problem Setup
Within the Supervised Fine-Tuning (SFT) framework, our VLN model based on NVILA 8B liu2024nvila , is learned by imitating expert demonstrations from the dataset, where each trajectory provides sequences of <(historical observations, current observation, instruction), expert action> pairs.
@@ 3.2 R2R-CoT-320k Dataset Construction
We present R2R-CoT-320k, the first VLN dataset annotated with CoT reasoning, tailored for the R2R-CE benchmark. We reconstruct step-wise navigation trajectories in the Habitat simulator habitat19iccv . Each sample in the dataset comprises the current view, the historical visual context, the corresponding instruction, and the ground-truth action. We employ Qwen-2.5-VL-72B bai2025qwen2 , one of the strongest publicly available VLMs, to generate detailed CoT for each navigation sample. For Pre-Think and Post-Think strategies, we format reasoning traces with <think></think> and <answer></answer> tags, following recent reasoning models guo2025deepseek . See Appendix A.2 for more details on our dataset.
@@ 3.3 Systematic Investigation on Reasoning Strategies for VLN
To investigate the impact of reasoning on VLN, we study and evaluate three distinct strategies for integrating Chain-of-Thought (CoT) reasoning during training and inference (Fig. 3(a)).
@@
The training loss for the VLN model $\pi_{\theta}$ above three strategies is:
@@
We isolate the impact of reasoning on overall navigation performance. In addition, we adjust the loss weight of the CoT part ($c_{t}^{*}$ in Equation 1) to make the model more focused on learning the action (Fig. 3(b)). Our key findings include:
(a) Illustration of different reasoning strategies.
(b) Comparision of success rate on Pre-Think, Post-Think, and No-Think Strategies.
Figure 3: Reasoning strategies and their performance. More detailed results are provided in the Section 4.4 and Appendix A.5.
Finding 1: Pre-Think and Post-Think perform significantly worse than No-Think. Despite involving explicit reasoning, both strategies lead to lower navigation success and efficiency, highlighting the unreliability of inference-time CoT in dynamic environments. Finding 2: A Balanced CoT Loss Weight Modestly Benefits Performance. This indicates that while inference-time CoT is problematic, the degree of emphasis on reasoning during training is a critical, strategy-specific factor in managing its effects on action learning.
We further analyze the severe performance degradation observed with Pre-Think and Post-Think strategies. During training, the model is only exposed to optimal sequences of states and actions in oracle trajectories. However, VLN environments are inherently complex, dynamic, and partially observable, increasing the likelihood of agents deviating from optimal, oracle-guided training trajectories. Consequently, when encountering non-oracle, out-of-distribution states during inference, the CoT reasoning generated by these strategies is susceptible to drift, potentially yielding inaccurate or hallucinated interpretations of the environment and instructions.
@@ 3.4 Aux-Think: Reasoning-Aware Co-Training Strategy
To address the challenges from the explicit CoT training to VLN, we propose Aux-Think, a novel reasoning strategy designed to enhance navigation performance without incurring inference-time problems. Aux-Think leverages reasoning exclusively during training through auxiliary tasks. We design two reasoning-based auxiliary tasks and one action-based primary task during training.
@@
Receding-Horizon Action Planning. We introduce Receding-Horizon Action Planning as our primary task. During training, the model predicts a sequence of the next $n$ actions $(a_{t},a_{t+1},...,a_{t+n-1})$ based on the instruction $I$, current observation $o_{t}$, and navigation history $\mathcal{O}_{t}$ for the sample at time step $t$ This setup encourages short-term forecasting while retaining reactivity to new observations. The training objective for each trajectory $\tau$ is defined as:
@@
For inference, we only activate the prediction of actions, and the model predicts the next $n$ actions and executes only the first one. This ensures fast, reactive navigation without reasoning overhead. We demonstrate its stabilizing effect in long-horizon trajectories through ablation studies (Table 5).
@@ 4.2 Implementation Details
Model training. We use NVILA-lite 8B liu2024nvila as the base pretrained model, which consists of a vision encoder (SigLIP tschannen2025siglip ), a projector, and an LLM (Qwen 2 bai2025qwen2 ). We use supervised finetuning (SFT) to train our VLN model from stage 2 of NVILA-lite, as it has finished visual language corpus pre-training. Our model is trained with 8 NVIDIA H20 GPUs for one epoch (around 60 hours), with a learning rate of 1e-5. In addition, we also try reinforcement fine-tuning, referring to Appendix A.4.
@@ 4.3 Comparison on VLN-CE Benchmarks
Open-Nav-Llama3.1qiao2024open Arxiv2024 $\checkmark$ 7.25 23 16 12.9 0K

### v4

# Yucheng Wang3,  Maiyue Chen3,  Kaihui Wang3,  Zhizhong Su3,  Deying Li1,  Zhaoxin Fan [^2] footnotemark: 2

@@ Abstract
Vision-Language Navigation (VLN) is a critical task for developing embodied agents that can follow natural language instructions to navigate in complex real-world environments. Recent advances driven by large pretrained models have significantly improved generalization and instruction grounding compared to traditional approaches. However, reasoning strategies in this task remain underexplored. Navigation is action-centric and long-horizon, while Chain-of-Thought (CoT) reasoning has mainly shown success in static tasks such as visual question answering. To address this gap, we conduct the first systematic evaluation of reasoning strategies, including No-Think (direct action prediction), Pre-Think (reasoning before action), and Post-Think (reasoning after action). Surprisingly, our findings reveal a Test-time Reasoning Collapse issue, where reasoning during testing degrades navigation accuracy, highlighting the challenges of integrating reasoning into embodied navigation. Based on this insight, we propose Aux-Think, a framework that trains models to internalize structured reasoning patterns via CoT supervision, while predicting actions directly without explicit reasoning at test time. To support this framework, we release R2R-CoT-320k, the first Chain-of-Thought annotated dataset for VLN. Extensive experiments show that Aux-Think substantially reduces training effort and achieves state-of-the-art performance on success rate.
@@ 1 Introduction
Vision-and-Language Navigation (VLN) wu2024vision ; anderson2018vision ; gu2022vision represents a groundbreaking step towards enabling robots to understand natural language instructions and navigate complex, unfamiliar environments. As a foundational capability for embodied AI systems, VLN bridges the gap between perception and action, empowering robots to seamlessly interact with the real world. In particular, Vision-Language Navigation in continuous environments (VLN-CE) hong2022bridging ; wang2023gridmm ; an2022bevbert has emerged as a critical research focus, pushing the boundaries of autonomy and adaptability in dynamic, real-world scenarios.
@@
Traditional Vision-and-Language Navigation methods often rely on waypoint predictors krantz2021waypoint ; hong2022bridging ; krantz2022sim or topology maps zubair2021sasra ; chen2021topological ; chen2022weakly ; wang2025mambavo but struggle with generalization and the sim-to-real gap. With the advancements in Large Language Models (LLMs) yang2024qwen2 ; touvron2023llama and Vision-Language Models (VLMs) bai2025qwen2 ; liu2024nvila ; lin2023video , recent studies zhang2024navid ; zhang2024uni ; cheng2024navila shift toward action prediction via supervised fine-tuning on paired videos and instructions. Despite these advancements, most efforts emphasize training strategies zhu2020vision ; zeng2024poliformer , data organization han2024roomtour3d , or model architecture zheng2024towards for VLN. Chain-of-Thought (CoT), which explicitly generates intermediate reasoning before producing final answers wei2022chain , has shown success in enhancing reasoning capabilities across various LLM- and VLM-driven tasks, like video understanding feng2025video and tool usage lu2025ui . However, its application to VLN remains unexplored.
@@
Motivated by this gap, we present the first systematic study of reasoning strategies in VLN, comparing three paradigms: (1) No-Think, which predicts actions without explicit reasoning; (2) Pre-Think zhang2024chain , which performs CoT before action selection; and (3) Post-Think li2025think , which reasons after action prediction. Our key findings reveal a phenomenon we term “Test-time Reasoning Collapse” (TRC): Introducing CoT via Pre-Think or Post-Think consistently harms navigation performance (Fig. 1(a)). CoT errors or hallucinations during testing result in incorrect actions, as shown in Fig. 3. We attribute this to a training–testing mismatch: CoT is only trained on optimal (oracle) trajectories, but for testing, agents often enter unfamiliar, off-distribution states where reasoning fails, leading to error accumulation along the trajectory and cascading navigation failures. Even with DAgger, coverage of off-distribution states is limited and CoT supervision remains biased toward the optimal region. This phenomenon highlights a fundamental limitation of multi-turn explicit reasoning in dynamic, partially observable environments, unlike the single-turn static tasks such as VQA or image captioning antol2015vqa ; wang2024videocot .
@@
To address the TRC issue, we propose Aux-Think, inspired by the dual-process theory of human learning evans2003two : during training, humans often rely on explicit reasoning to understand principles, but during execution, they focus on actions without consciously recalling those principles, much like a driver who no longer recites traffic rules while driving. Similarly, Aux-Think uses CoT as an auxiliary signal during training to guide the model in internalizing reasoning patterns. At testing time, the model no longer generates explicit reasoning, but instead directly predicts actions based on the internalized reasoning learned during training. This separation between learning and execution improves decision focus, reduces testing overhead and hallucinations, and leads to more accurate and stable navigation.
@@
Specifically, in Aux-Think, the generation of CoT reasoning and navigation actions is decoupled into two distinct tasks during training: (1) generating navigation actions based on stepwise observations and language instructions as the primary task, and (2) generating the reasoning process for each step as an auxiliary task. By leveraging prompts to switch between these tasks, we independently supervise navigation actions and reasoning processes, effectively avoiding the negative interference caused by jointly training both tasks. During testing, Aux-Think directly predicts actions without intermediate reasoning, thereby eliminating the risk of errors introduced by CoT hallucinations.
@@
  • New Finding: We conduct a systematic comparison of reasoning strategies in VLN and reveal that test-time reasoning, including Pre-Think and Post-Think, consistently underperforms direct action prediction (No-Think), termed the TRC issue. To our knowledge, this is the first exploration of CoT strategies on the VLN-CE task.
  • New Method: We propose Aux-Think, a novel training paradigm that uses CoT as auxiliary supervision while maintaining No-Think testing, achieving superior performance over other reasoning methods. Aux-Think pioneers a new perspective on CoT utilization and achieves the best performance on the navigation success rate.
@@
Figure 2: Illustration of Aux-Think and other reasoning strategies. Unlike No-Think, Pre-Think, and Post-Think, our Aux-Think introduces auxiliary CoT- and instruction-based reasoning during training while maintaining efficient action planning at testing.
Figure 3: Illustration of CoT and Action Prediction Results Using Different Reasoning Strategies. Pre-Think generates incorrect actions (e.g., “move forward 75cm”) due to flawed CoT reasoning, such as “we have not crossed the room,” leading to significant trajectory deviation. Post-Think, which builds on Pre-Think’s output, inherits similar reasoning errors (e.g., “no obvious door or archway”) and makes the same wrong prediction. In contrast, Aux-Think correctly predicts “turn right 15 degrees” and follows a trajectory aligned with the ground truth. While Aux-Think does not rely on CoT during testing, it can optionally produce CoT via prompt switching—yet its action prediction remains accurate even when the generated CoT is of moderate quality. This highlights Aux-Think’s robustness to imperfect reasoning and its superior reliability in action prediction.
@@ 2.2 VLN with Large Pretrained Models
Recent advancements have seen the integration of large pre-trained models liu2024nvila ; bai2025qwen2 ; yang2024qwen2 ; wang2025monodream , into VLN tasks. Early explorations of LLM in the VLN field usually use off-the-shelf large language models to select landmarks or waypoints in a zero-shot manner long2024discuss ; long2024instructnav ; shah2023lm ; zhou2024navgpt . Recent works have focused on fine-tuning the VLM to obtain the navigational Vision-Language-Action model. Notably, Poliformer zeng2024poliformer and NaVid zhang2024navid introduce a video-based monocular VLN, demonstrating navigation capabilities using monocular RGB videos without maps or depth input. Uni-NaVid zhang2024uni unifies various navigation tasks, including VLN, ObjectNav chaplot2020object , Embodied Question Answering das2018embodied , and Human-following islam2019person ; puig2023habitat , into a single model trained on a diverse dataset. NaVILA cheng2024navila further extends this approach by integrating VLN with legged robot locomotion skills in complex environments.
While these models have improved the alignment among visual understanding, language instructions, and navigation actions, they predominantly employ No-Think testing strategies, lacking reasoning mechanisms. Moreover, their performance gains often stem from leveraging extensive datasets, whereas our approach focuses on exploring reasoning strategies.
@@ 2.3 Reasoning Models
In our work, we conduct the first systematic comparison of Pre-Think, Post-Think, and No-Think reasoning strategies for VLN. Based on our findings, we propose Aux-Think, a novel framework that leverages CoT reasoning as auxiliary supervision during training while maintaining No-Think testing, thereby enhancing data efficiency and performance in VLN.
@@ 3.1 Problem Setup
Within the Supervised Fine-Tuning (SFT) framework, our VLN model based on NVILA 8B liu2024nvila , is learned by imitating expert demonstrations from the dataset, where each trajectory provides sequences of <navigation context, expert action> pairs, with navigation context denoting the combination of historical observations, the current observation, and the instruction.
@@ 3.2 R2R-CoT-320k Dataset Construction
We present R2R-CoT-320k, the first VLN dataset annotated with CoT reasoning, tailored for the R2R-CE benchmark. We reconstruct step-wise navigation trajectories in the Habitat simulator habitat19iccv . Each sample in the dataset comprises the current view, the historical visual context, the corresponding instruction, and the ground-truth action. We employ Qwen-2.5-VL-72B bai2025qwen2 , one of the strongest publicly available VLMs, to generate detailed CoT for each navigation sample (Fig. 4). For Pre-Think and Post-Think strategies, we format reasoning traces with <think></think> and <answer></answer> tags, following recent reasoning models guo2025deepseek . See Appendix A.2 for more details on our dataset.
Figure 4: The annotation pipeline of our R2R-CoT-320k dataset.
@@ 3.3 Systematic Investigation on Reasoning Strategies for VLN
To investigate the impact of reasoning on VLN, we study and evaluate three distinct strategies for integrating Chain-of-Thought (CoT) reasoning during training and testing.
Figure 5: Comparision of success rate on Pre-Think, Post-Think, and No-Think.
@@
The training loss for the VLN model $\pi_{\theta}$ with above three strategies is:
@@
We isolate the impact of reasoning on overall navigation performance. In addition, we adjust the loss weight of the CoT part ($c_{t}^{*}$ in Equation 1) to make the model more focused on learning the action (Fig. 5). Our key findings include:
Finding 1: Pre-Think and Post-Think perform significantly worse than No-Think. Despite involving explicit reasoning, both strategies lead to lower navigation success and efficiency, highlighting the unreliability of test-time CoT in dynamic environments. Finding 2: Careful Balancing of Explicit CoT Loss Weight Improves Performance. By carefully tuning the explicit CoT weight for Pre-Think and Post-Think, we find that balanced supervision yields slight gains, indicating that training emphasis on reasoning is a key, strategy-dependent factor despite test-time unreliability.
We further analyze the severe performance degradation observed with Pre-Think and Post-Think strategies. During training, the model is only exposed to optimal sequences of states and actions in oracle trajectories. However, VLN environments are inherently complex, dynamic, and partially observable, increasing the likelihood of agents deviating from optimal, oracle-guided training trajectories. Consequently, when encountering non-oracle, out-of-distribution states during testing, the CoT reasoning generated by these strategies is susceptible to drift, potentially yielding inaccurate or hallucinated interpretations of the environment and instructions.
@@ 3.4 Aux-Think: Reasoning-Aware Co-Training Strategy
To address the challenges from the explicit CoT training to VLN, we propose Aux-Think, a novel reasoning strategy designed to enhance navigation performance without incurring test-time problems. Aux-Think leverages reasoning exclusively during training through auxiliary tasks. We design two reasoning-based auxiliary tasks and one action-based primary task during training.
@@
Receding-Horizon Action Planning. We introduce Receding-Horizon Action Planning as our primary task. During training, the model predicts a sequence of the next $n$ actions $(a_{t},a_{t+1},...,a_{t+n-1})$ based on the instruction $I$, current observation $o_{t}$, and navigation history $\mathcal{O}_{t}$ for the sample at time step $t$ . This setup encourages short-term forecasting while retaining reactivity to new observations. The training objective for each trajectory $\tau$ is defined as:
@@
For testing, we only activate the prediction of actions, and the model predicts the next $n$ actions and executes only the first one. This ensures fast, reactive navigation without reasoning overhead. We demonstrate its stabilizing effect in long-horizon trajectories through ablation studies (Table 5).
@@ 4.2 Implementation Details
Model training. We use NVILA-lite 8B liu2024nvila as the base pretrained model, which consists of a vision encoder (SigLIP tschannen2025siglip ), a projector, and an LLM (Qwen 2 bai2025qwen2 ). We use supervised finetuning (SFT) to train our VLN model from stage 2 of NVILA-lite, as it has finished visual language corpus pre-training. Our model is trained with 8 NVIDIA H20 GPUs for one epoch (around 60 hours), with a learning rate of 1e-5.